# Neuro-Symbolic Agents with Modal Logic for Autonomous Diagnostics

### v1

@@ 1 Introduction
Simply making an environment look more realistic or adding more complexity is not enough to create the next generation of truly capable AI agents. For these agents to operate independently and reliably, especially in high-risk areas like industrial process control, automated scientific research and autonomous accelerator control sulc2024towards we need to understand how they think, be able to formally check their reasoning, and ensure their decision-making is firmly based on the core principles of the field they work in. Language Models (LMs) are powerful tools for processing and generating language, but they sometimes produce answers that sound convincing yet are factually wrong or logically flawed. This problem, often called hallucination zhang2023siren, makes it risky to use them in critical systems where safety and reliability are essential ji2023survey ; zhang2025siren .
@@ 2.1 Kripke Models as Agent Belief States
  • $W$ is a set of possible worlds. For a diagnostic agent, a ”world” represents a complete and plausible state of the entire environment. For instance, a world $w_{0}$ might represent a state where the system is operating nominally, while another world $w_{1}$ could represent a state where a specific cooling fault has occurred.
  • $R$ is an accessibility relation defined on $W$. A relation $(w_{i},w_{j})\in R$ signifies that if the agent currently believes it is in world $w_{i}$, it considers world $w_{j}$ to be a possible future or alternative state, thereby defining the agent’s uncertainty and its internal model of state transitions.
  • $V$ is a valuation function that assigns a set of true atomic propositions to each world $w\in W$. For example, the valuation of world $w_{1}$ might be $V(w_{1})=\{\text{pressure\_low},\text{cooling\_fault\_reported}\}$.
@@ 3 A Multi-Agent Architecture for Diagnostics
  • Component Monitoring Agents: These serve as the observers, with each agent being responsible for a specific subsystem (e.g. cooling or radio-frequency power). These agents monitor a small, dedicated set of signals and utilize their local neuro-symbolic loop to form initial, localized hypotheses when anomalies are detected.
  • Hierarchical Reasoning Agent: This agent controls the monitors and does not observe environmental signals directly. Instead, its inputs are the structured reports and belief models from the various component agents. Its primary role is to synthesize these disparate reports into a single, coherent causal theory for the entire system, and it is endowed with the most comprehensive set of expert rules.
  • Physical Knowledge Agent: This unique agent acts as a specialized repository of factual information. It holds no beliefs about the dynamic state of the environment; its knowledge base is a static, factual representation of the system’s physical topology. Its sole purpose is to answer specific queries from the Reasoning Agent to verify the physical plausibility of a given causal theory.
@@ 7 Discussion and Future Work
Our work demonstrates a promising and viable path for building more reliable and trustworthy autonomous agents. By treating the LM not as an infallible oracle but as a powerful hypothesis generator whose outputs are subject to rigorous symbolic verification, we can effectively harness its strengths while simultaneously mitigating its known weaknesses. The use of modal logic and Kripke models provides a formal semantics for agent belief, which is a critical and necessary step towards creating agents that can transparently explain their reasoning and provide formal guarantees about their behavior. Further exploration into Dynamic Epistemic Logic could enhance the model’s ability to reason about how agents’ knowledge changes as a result of new information or actions van2007dynamic .
This framework opens up several exciting directions for future research. The current process of translating LM outputs into formal propositions is straightforward; developing more complex semantic parsing techniques, possibly using methodologies from the field of computational linguistics, could allow for the expression of richer and more nuanced hypotheses. Future work could explore advanced methods for learning these logical constraints directly from data or through guided interaction with human experts, potentially leading to systems that can formalize their own operational theories over time, a concept related to theory of mind in AI devin2024learning . Finally, applying this architecture to an online reinforcement learning setting, where an agent’s actions have tangible consequences and its belief models are continually updated based on rich environmental feedback, represents the next logical step toward achieving true end-to-end autonomy.
@@ A.2 Simplifications and Justifications
Appendix B Grounding Semantic Hypotheses into Logical Propositions
The process is not a ”magic” translation but a constrained, multi-stage engineering solution:
  1. 1.
    Predefined and Finite Vocabulary: The symbolic part of our system operates on a small, predefined set of atomic propositions (e.g., ‘cooling_fault_reported,‘klystron_fault_reported,‘rf_power_low). These propositions are explicitly defined in each agent’s initial Kripke model and constitute the complete logical vocabulary for the diagnostic task. The system does not need to invent or parse novel logical terms.
  2. 2.
    Structured Prompting and Classification: We do not ask the Language Model (LM) to generate a logical formula from scratch. Instead, we use structured prompting to constrain the LM’s task to one of classification. When an anomaly is detected, the LM is given a prompt that explicitly asks it to categorize the fault into one of the predefined system types. The prompt for the monitoring agents includes the following instruction:
    ”Your response MUST be a JSON object with one key: ’suspected_system’, which should be one of the following: ’Cooling’, ’Power’, ’Vacuum’, ’Klystron’…”
    This turns the LM into a powerful semantic classifier rather than an open-ended generator. The LM’s role is to use its world knowledge to map the semantics of an anomaly report (e.g., ”high temperature”) to the most likely faulty subsystem (”Cooling”) from a fixed list of options.
  3. 3.
    Deterministic Mapping: The final translation step is a simple, deterministic mapping handled by the agent’s code. The agent receives the structured JSON output from the LM and uses a hard-coded dictionary or switch-case statement to map the string output to the corresponding formal proposition. For example:

### v2

# Code available at https://github.com/sulcantonin/neuro-symbolic-diagnostics.git

@@ 1 Introduction
Simply making an environment look more realistic or adding more complexity is not enough to create the next generation of truly capable AI agents. For these agents to operate independently and reliably, especially in high-risk areas like industrial process control, automated scientific research and autonomous accelerator control sulc2024towards we need to understand how they think, be able to formally check their reasoning, and ensure their decision-making is firmly based on the core principles of the field they work in. Language Models (LMs) are powerful tools for processing and generating language, but they sometimes produce answers that sound convincing yet are factually wrong or logically flawed. This problem, often called hallucination, makes it risky to use them in critical systems where safety and reliability are essential ji2023survey ; zhang2025siren .
@@ 2.1 Kripke Models as Agent Belief States
First $W$ stands for a set of possible worlds. For a diagnostic agent, a ”world” represents a complete and plausible state of the entire environment. For instance, a world $w_{0}$ might represent a state where the system is operating nominally, while another world $w_{1}$ could represent a state where a specific cooling fault has occurred. Second element is $R$, which denotes an accessibility relation defined on $W$. A relation $(w_{i},w_{j})\in R$ signifies that if the agent currently believes it is in world $w_{i}$, it considers world $w_{j}$ to be a possible future or alternative state, thereby defining the agent’s uncertainty and its internal model of state transitions. Lastly, $V$ is a valuation function that assigns a set of true atomic propositions to each world $w\in W$. For example, the valuation of world $w_{1}$ might be $V(w_{1})=\{\text{pressure\_low},\text{cooling\_fault\_reported}\}$.
@@ 2.2 The Neuro-Symbolic Loop
2.3 Grounding Semantic Hypotheses into Logical Propositions
The symbolic part of our system operates on a small, predefined set of atomic propositions (e.g., ‘cooling_fault_reported, ‘klystron_fault_reported,‘rf_power_low). These propositions are explicitly defined in each agent’s initial Kripke model and constitute the complete logical vocabulary for the diagnostic task. The system does not need to invent or parse novel logical terms.
The process is a constrained, multi-stage engineering solution. First, we do not ask the LM to generate a logical formula from scratch. Instead, we use structured prompting to constrain the LM’s task to one of classification. When an anomaly is detected, the LM is given a prompt that explicitly asks it to categorize the fault into one of the predefined system types. The prompt for the monitoring agents includes the following instruction:
Your response MUST be a JSON object with one key: ’suspected_system’, which should be one of the following: ’Cooling’, ’Power’, ’Vacuum’, ’Klystron’…
This structured prompting turns the LM into a powerful semantic classifier rather than an open-ended generator with rich internal knowledge about accelerators. The LM’s role is to use its world knowledge to map the semantics of an anomaly report (e.g., ”high temperature”) to the most likely faulty subsystem (”Cooling”) from a fixed list of options.
The final translation step is a simple, deterministic mapping handled by the agent’s code. The agent receives the structured JSON output from the LM and uses a hard-coded dictionary or switch-case statement to map the string output to the corresponding formal proposition. For example:
This allows integration of the LMs with their rich knowledge for diagnosis of our complex scenario, while their outputs are guard-railed to stick to an explainable scenario.
@@ 3 A Multi-Agent Architecture for Diagnostics
Component Monitoring Agents:
These serve as the observers, with each agent being responsible for a specific subsystem (e.g. cooling or radio-frequency power). These agents monitor a small, dedicated set of signals and utilize their local neuro-symbolic loop to form initial, localized hypotheses when anomalies are detected.
Hierarchical Reasoning Agent:
This agent controls the monitors and does not observe environmental signals directly. Instead, its inputs are the structured reports and belief models from the various component agents. Its primary role is to synthesize these disparate reports into a single, coherent causal theory for the entire system, and it is endowed with the most comprehensive set of expert rules.
Physical Knowledge Agent:
This unique agent acts as a specialized repository of factual information. It holds no beliefs about the dynamic state of the environment; its knowledge base is a static, factual representation of the system’s physical topology. Its sole purpose is to answer specific queries from the Reasoning Agent to verify the physical plausibility of a given causal theory.
@@ 7 Discussion and Future Work
Our work demonstrates a promising and viable path for building more reliable and trustworthy autonomous agents for a fast diagnosis at accelerator controls. By treating the LM not as an infallible oracle but as a powerful hypothesis generator whose outputs are subject to rigorous symbolic verification, we can effectively harness its strengths while simultaneously mitigating its known weaknesses. The use of modal logic and Kripke models provides a formal semantics for agent belief, which is a critical and necessary step towards creating agents that can transparently explain their reasoning and provide formal guarantees about their behavior. Further exploration into Dynamic Epistemic Logic could enhance the model’s ability to reason about how agents’ knowledge changes as a result of new information or actions van2007dynamic .
This framework opens up several exciting directions for future research. The current process of translating LM outputs into formal propositions is straightforward; developing more complex semantic parsing techniques, possibly using methodologies from the field of computational linguistics, could allow for the expression of richer and more nuanced hypotheses. Future work could explore advanced methods for learning these logical constraints directly from data or through guided interaction with human experts, potentially leading to systems that can formalize their own operational theories over time, a concept related to theory of mind in AI runge2020discovering . Finally, applying this architecture to an online reinforcement learning setting, where an agent’s actions have tangible consequences and its belief models are continually updated based on rich environmental feedback, represents the next logical step toward achieving true end-to-end autonomy.